# Graph-Structured Topic Modeling for Documents with Spatial or Covariate Dependencies

### v1

@@ B.3 Analysis of the Estimation of $A$
$\begin{split}\|X-\widehat{W}\widehat{A}\|_{F}^{2}&\leq\|X-\widehat{W}P^{-1}A\|_{F}^{2}\\ \|WA+Z-\widehat{W}\widehat{A}\|_{F}^{2}&\leq\|WA+Z-\widehat{W}P^{-1}A\|_{F}^{2}\\ \|(W-\widehat{W}P^{-1})A+\widehat{W}(P^{-1}A-\widehat{A})+Z\|_{F}^{2}&\leq\|(W-\widehat{W}P^{-1})A+Z\|_{F}^{2}\\ \|\widehat{W}(P^{-1}A-\widehat{A})\|_{F}^{2}&\leq 2\langle\widehat{W}(\widehat{A}-P^{-1}A),(W-\widehat{W}P^{-1})A+Z\rangle\\ &=2\langle\widehat{W}(\widehat{A}-P^{-1}A),(W-\widehat{W}P^{-1})A\rangle+2\langle\widehat{W}(\widehat{A}-P^{-1}A),Z\rangle\\ &\leq 2\|\widehat{W}(\widehat{A}-P^{-1}A)\|_{F}\|(W-\widehat{W}P^{-1})A\|_{F}\\ &+2\|\widehat{W}(\widehat{A}-P^{-1}A)\|_{F}\max_{U\in\mathbb{R}^{n\times p}:\|U\|_{F}=1}\langle U,Z\rangle\\ \|\widehat{W}(P^{-1}A-\widehat{A})\|_{F}&\leq 2\|(W-\widehat{W}P^{-1})A\|_{F}+C_{2}\sqrt{\frac{\log(n)}{N}}\\ &\leq 2\lambda_{1}(A)\|(W-\widehat{W}P^{-1})\|_{F}+C_{2}\sqrt{\frac{\log(n)}{N}}\\ \|P^{-1}A-\widehat{A}\|_{F}&\leq 2\|(W-\widehat{W}P^{-1})A\|_{F}+C_{2}\sqrt{\frac{\log(n)}{N}}\\ &\leq 2\frac{\lambda_{1}(A)}{\lambda_{\min}(\widehat{W})}\|W-\widehat{W}P^{-1}\|_{F}+C_{2}\sqrt{\frac{\log(n)}{N}}\\ &\leq\frac{\lambda_{1}(A)\|W-\widehat{W}P^{-1}\|_{F}}{\lambda_{K}(W)-\|W-\widehat{W}P^{-1}\|_{op}}+2C_{2}\sqrt{\frac{\log(n)}{N}}\qquad(*)\\ &\leq\frac{\lambda_{1}(A)}{\lambda_{K}(W)/\|W-\widehat{W}P^{-1}\|_{F}-1}+2C_{2}\sqrt{\frac{\log(n)}{N}}\\ &\leq C_{1}\frac{\lambda_{1}(A)}{\lambda_{K}(W)}\|W-\widehat{W}P^{-1}\|_{F}+2C_{2}\sqrt{\frac{\log(n)}{N}}\end{split}$
@@
Now, assume that $N$ is large enough so that (12) holds. Combining Lemma 5 and Theorem 3, $\|W-\widehat{W}P^{-1}\|_{F}$ becomes small enough so that $\|W-\widehat{W}P^{-1}\|_{F}\leq\frac{2}{c}<1\leq\lambda_{K}(W)$.
@@
Lastly, we use the fact, $\lambda_{1}(A)\leq\|A\|_{F}\leq\sqrt{K}$, and $\lambda_{K}(W)\geq 1$ from Lemma 5 and the result of Theorem 3 to get the final bound of $A$,
@@
B.4 Comparison with One-step Graph-Aligned denoising
We also propose a fast one-step graph-aligned denoising of the matrix $X$ that could be an alternative of the iterative graph-aligned SVD in Step 1 of 2.3. We denoise the frequency matrix $X$ by the following optimization problem,
A SVD on the denoised matrix $\widehat{M}$ yields estimates of the singular values $U$ and $V$ (Algorithm 3). Through extensive experiments with synthetic data detailed in Section 4, we find that one-step graph-aligned denoising provides more accurate estimates than pLSI but still falls short compared to the iterative graph-aligned denoising (GpLSI). We provide a theoretical upper bound on its error as well as its comparison to the error of pLSI where there is no graph-aligned denoising.
Algorithm 3 One-step Graph-aligned denoising
1:Input: Observation $X$, incidence matrix $\Gamma$
2:Output: Denoised singular vectors $\widehat{U}$ and $\widehat{V}$.
We begin by analyzing the one-step graph-aligned denoising, as proposed in Algorithm 3. We begin by reminding the reader that, in our proposed setting, the observed word frequencies in each document are assumed to follow a “signal + noise” model, as per (2):
$X=M+Z$
where the true probability $M$ is assumed to admit the following SVD decomposition:
$M=\mathbb{E}[X]=U\Lambda V^{\top}.$
Theorem 5 (Upper bound on the one-step denoising error)
Let Assumptions 1 to 5 hold. Denote $\rho(\Gamma),s,n_{\mathcal{C}}$, and $n_{\mathcal{C}_{\min}}$ as (3)-(5) and assume $\max(K,p)\leq n$ and $\sqrt{K}\leq p$. Let $\widehat{U}$ and $\widehat{V}$ be given as estimators obtained from Algorithm 3. Then, there exists a constant $C>0$, such that with probability at least $1-o(n^{-1})$,
where $\lambda_{K}(M)$ denotes the $K^{th}$ singular value of $M$.
Proof
Let $\widehat{M}$ be the solution of (51):
Let $\Delta=\widehat{M}-M$, and $Z=X-M$. By the basic inequality, we have:
$\forall x\in\mathbb{R}^{n},\qquad x=\Pi x+\Gamma^{\dagger}\Gamma x$
B.4.0.1 Concentration of (A).
By Cauchy-Schwarz, it is immediate to see that:
$\langle\Pi Z,\Pi\Delta\rangle\leq\|\Pi Z\|_{F}\|\Pi\Delta\|_{F}.$
By Lemma 12, with probability at least $1-o(n^{-1})$:
$(A)\leq 2\sqrt{C_{1}Kn_{C}{\frac{\log(n)}{N}}}\|\Delta\|_{F}$
B.4.0.2 Concentration of (B).
We have:
Then, on $\mathcal{A}$, we have:
B.4.0.3 Concentration
We thus have:
$\begin{split}\|\Delta\|_{F}^{2}&\leq 4\sqrt{C_{1}Kn_{C}{\frac{\log(n)}{N}}}\|\Delta\|_{F}+\frac{3\rho}{2}\|(\Gamma\Delta)_{S\cdot}\|_{21}\\ &\leq 4\sqrt{C_{1}Kn_{C}{\frac{\log(n)}{N}}}\|\Delta\|_{F}+\frac{3\rho}{2}\sqrt{s}\|\Gamma\Delta\|_{F}\\ &\leq 4\sqrt{C_{1}Kn_{C}{\frac{\log(n)}{N}}}\|\Delta\|_{F}+\frac{3\rho}{2}\sqrt{s}\sqrt{\lambda_{\max}(\Gamma)}\|\Delta\|_{F}\\ \|\Delta\|_{F}&\leq 4\sqrt{C_{1}Kn_{C}{\frac{\log(n)}{N}}}+\frac{3\rho}{2}\sqrt{s}\sqrt{\lambda_{\max}(\Gamma)}\\ \|\Delta\|_{F}&\leq 4\sqrt{C_{1}Kn_{C}{\frac{\log(n)}{N}}}+6\rho(\Gamma)C_{2}\sqrt{\frac{K\log(n)}{N}}\sqrt{s}\sqrt{\lambda_{\max}(\Gamma)}\\ &\leq C\left(\sqrt{n_{C}}+\rho(\Gamma)\sqrt{s}\sqrt{\lambda_{\max}(\Gamma)}\right)\sqrt{\frac{K\log(n)}{N}}\end{split}$
[…]
$\max\{\|\sin\Theta(U,\widehat{U})\|_{F},\|\sin\Theta(V,\widehat{V})\|_{F}\}\leq\frac{C}{\lambda_{K}(M)}\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{C}}+\rho(\Gamma)\sqrt{s}\sqrt{\lambda_{\max}(\Gamma)}\right)$
This concludes our proof.
We observe first that the error bound for one-step graph-aligned denoising has better rate than the one with no regularization, provided in (14). We also note that the rate of one-step denoising and GpLSI is equivalent up to a constant. Although the dependency of the error on parameters $n,p,K,$ and $N$ is the same for both methods, our empirical studies in Section 3.4 reveal that GpLSI still achieves lower errors compared to one-step denoising.
Appendix C Auxiliary Lemmas
Let matrices $X,M,Z,W,\text{ and }A$ be defined as in Equation (2). In this section, we provide inequalities on the singular values of the unobserved quantities $W,M,\text{ and }H$, perturbation bounds for singular spaces, as well as concentration bounds on noise terms, which are useful for proving our main results in Section 3.
C.1 Inequalities on Unobserved Quantities
Lemma 4
For the matrix $M$ (Equation 2), the following inequalities hold:
@@ C.1 Inequalities on Unobserved Quantities
Proof  We observe that for each $j\in[p]$, the variational characterization of the smallest eigenvalue of the matrix $M^{\top}M/n$ yields:
@@
Lemma 5 (Lemma 6 from the supplemental material of Klopp et al. (2021))
Let Assumption 2 be satisfied. For the matrices $W$, $H$, $\widehat{H}$ defined in Equations (6) and (7), we have
@@
C.2 Matrix Perturbation Bounds
In this section, we provide rate-optimal bounds for the left and right singular subspaces. While the original Wedin’s perturbation bound (Wedin, 1972) treats the singular subspaces symmetrically, work by Cai and Zhang (2018) provides sharper bounds for each subspace individually. This refinement is particularly relevant in our setting where an additional denoising step of the left singular subspace leads to different perturbation behaviors of left and right singular subspaces as iterations progress.
@@ C.2 Matrix Perturbation Bounds
Let $X=M+Z$ be the perturbed version of $M$ with the perturbation matrix $Z$. We can write the SVD of $X$ as:
@@
Lemma 6 (Adapted from Theorem 1 of Cai and Zhang (2018))
Let $M$, $X$, and $Z$ be given as in Equations (57)-(59). Then:
@@
Proof  This result is a simplified version of the original theorem under the setting $rank(M)=K$.
C.3 Concentration Bounds
We first introduce the general Bernstein inequality and its variant which will be used for proving high probability bounds for noise terms in Section C.3.2.
C.3.1 General Inequalities
Lemma 7 (Bernstein inequality (Corollary 2.11, Boucheron et al. (2013)))
@@ C.3.1 General Inequalities
Lemma 8 (Bernstein inequality for bounded variables (Theorem 2.8.4, Vershynin (2018)))
@@
C.3.2 Technical Lemmas
Lemma 9 (Concentration of the cross-terms $Z_{i}^{T}Z_{j}$)
Let Assumptions 1-4 hold. With probability at least $1-o(n^{-1})$:
@@ C.3.2 Technical Lemmas
Proof  The proof is a re-adaptation of Lemma C.4 in Tran et al. (2023) for any word $j$
Similar to the analysis of Ke and Wang (2017), we rewrite each row $X_{i\cdot}$ as a sum over $N$ word entries $T_{im}\in\mathbb{R}^{p}$, where $T_{im}$ denotes the $m^{th}$ word in document $i$, encoded as a one-hot vector:
@@
In the previous expression, under the pLSI model (Equation 1), the $\{T_{im}\}_{m=1}^{N}$ are i.i.d. samples from a multinomial distribution with parameter $M_{i\cdot}$.
@@
so $Q$ is a (double) summation of mutually independent variables. We have $|Q_{im}|\leq 1$, $\mathbb{E}(Q_{im})=0$ and $\mathbb{E}(Q^{2}_{im})\leq{h}_{j}{h}_{l}$ where $\forall j\in[p],h_{j}=\sum_{k=1}^{K}A_{kj}$.. The rest of the proof for $V_{2}$ is similar to the standard proof for the usual Bernstein’s inequality.
@@
Lemma 10 (Concentration of the covariance matrix $X^{\top}X$)
@@
[…]
@@
since we assume that $p\ll n$. Therefore, with probability at least $o(n^{-1})$:
@@
Noting that $\lambda_{1}(M)\leq\sqrt{n}$ (Lemma 4), by Lemma 13, with probability at least $1-o(n^{-1})$:
@@
Lemma 11 (Concentration of $\|(\Gamma^{\dagger}Z)_{e\cdot}\|_{2},e\in\mathcal{E}$)
@@
    Each $\eta_{im}$ verifies Bernstein’s condition (Equation (62)): We have:
@@
    Each of the variance $\text{Var}(S_{m})=\sum_{i=1}^{n}\text{Var}(\eta_{im})$ is also bounded:
@@
with $c_{0}<2$. Therefore, $\mathbb{P}[\exists e\in\mathcal{E}:\quad\|(\Gamma^{\dagger})^{\top}Z)_{e\cdot}\|^{2}\geq C^{*}\rho^{2}(\Gamma){\frac{K\log(n)}{N}}]=o(\frac{1}{n})$ for a choice of $C^{*}$ sufficiently large.
Lemma 12 (Concentration of $\|\Pi Z\|_{F}$)
@@
Proof  We remind the reader that letting $\mathcal{C}_{j}$ denote the $j^{th}$ connected component of the graph $\mathcal{G}$ and $n_{\mathcal{C}_{l}}=|\mathcal{C}_{l}|$ its cardinality, $\Pi$ can be arranged as a block diagonal form where each block represents a connected component, $\Pi_{[\mathcal{C}_{l}]}=\frac{1}{n_{\mathcal{C}_{l}}}\mathbf{1}_{\mathcal{C}_{l}}\mathbf{1}_{\mathcal{C}_{l}}^{T}$. Since the components $\mathcal{C}_{l}$ are all disjoint, $\|\Pi Z\|_{F}$ can be further decomposed as:
$\begin{split}\|\Pi Z\|_{F}^{2}&\leq\sum_{l=1}^{n_{C}}\|\frac{1}{n_{\mathcal{C}_{l}}}\mathbf{1}_{\mathcal{C}_{l}}\mathbf{1}_{\mathcal{C}_{l}}^{T}Z_{[\mathcal{C}_{l}]}\|_{F}^{2}\\ &=\sum_{l=1}^{n_{C}}n_{\mathcal{C}_{l}}\|\frac{1}{n_{\mathcal{C}_{l}}}\mathbf{1}_{\mathcal{C}_{l}}^{T}Z_{[\mathcal{C}_{l}]}\|_{2}^{2}\end{split}$
@@
In the previous expression, under the pLSI model, (1), the $\{T_{im}\}_{m=1}^{N}$ are i.i.d. samples from a multinomial distribution with parameter $M_{i\cdot}$. Thus, for each word $j$ and each connected component $\mathcal{C}_{l}$:
$\begin{split}\frac{1}{n_{\mathcal{C}_{l}}}\mathbf{1}_{\mathcal{C}_{l}}^{T}Z_{[\mathcal{C}_{l}]j}&=\frac{1}{n_{\mathcal{C}_{l}}N}\sum_{i\in\mathcal{C}_{l}}\sum_{m=1}^{M}(T_{im}(j)-\mathbb{E}[T_{im}(j)]).\end{split}$
@@
$\begin{split}&\mathbb{P}\left[\exists j\in[p],\exists l\in[n_{C}]:\quad\frac{1}{n_{\mathcal{C}_{l}}N}|\sum_{i\in\mathcal{C}_{l}}\sum_{m=1}^{N}S^{(j)}_{im}|>\sqrt{C^{*}\frac{h_{j}}{n_{\mathcal{C}_{l}}N}\log(n)}\right]\\ &\leq 2pn_{C}\exp\{-{C^{*}\log(n)}\}\\ &=\exp\{\log(2)+\log(p)+\log(n_{C})-C^{*}\log(n)\}\\ &\leq\exp\{-C\log(n)\},\end{split}$
where the last line holds for some large enough constant $\tilde{C}<C^{*}$ such that $n^{\tilde{C}}\geq 2pn_{C}$ and $C=C^{*}-\tilde{C}$. As a consequence of Assumption 5, we know that $p\ll n$ (see remark 2) and under the graph-aligned setting, $n_{\mathcal{C}}\ll n$. Thus with probability $1-o(n^{-1})$, for all $j\in[p]$ and all $l\in[n_{C}]$:
$\begin{split}\|\frac{1}{n_{l}}\mathbf{1}_{\mathcal{C}_{l}}^{T}Z_{[n_{C}]}\|_{2}^{2}&\leq\sum_{j\in[p]}C^{*}\frac{h_{j}}{n_{\mathcal{C}_{l}}N}\log(n)=C^{*}\frac{K}{n_{\mathcal{C}_{l}}N}\log(n).\end{split}$
@@
Lemma 13 (Concentration of $\|U^{\top}Z\|_{F}$)
Let Assumptions 1-5 hold. Let $U\in\mathbb{R}^{n\times r}$ denote a projection matrix: $U^{T}U=I_{r}$, with $r$ a term that does not grow with $n$ or $p$ and $r\leq n$, and let $Z$ denote some centered multinomial noise as in Equation 2. Then with probability at least $1-o(n^{-1})$:
@@
Proof  Let $\tilde{Z}=U^{\top}Z.$ We have:
@@
Thus, ${\tilde{Z}}_{kj}$ is a sum of $nN$ centered independent variables.
@@
Since we assume that $pr\ll n$, the result follows.
Lemma 14 (Concentration of $\|\Pi ZV\|_{F}$)
Let Assumptions 1-5 hold. Let $V$ be a orthogonal matrix: $V\in\mathbb{R}^{p\times K},V^{\top}V=I_{K}$. Let $\Pi$ denote the projection matrix unto $\text{Ker}(\Gamma^{\dagger}\Gamma)$, such that $I_{n}=\Pi\oplus^{\perp}\Gamma^{\dagger}\Gamma$. With probability at least $o(n^{-1})$:
@@
Proof  We follow the same procedure as the proof of Lemma  12. Letting $\mathcal{C}_{j}$ denote the $j^{th}$ connected component of the graph $\mathcal{G}$ and $n_{\mathcal{C}_{l}}=|\mathcal{C}_{l}|$ its cardinality, $\|\Pi\tilde{Z}\|_{F}$ can be decomposed as:
@@
$\begin{split}\frac{1}{n_{\mathcal{C}_{l}}}\mathbf{1}_{\mathcal{C}_{l}}^{T}\tilde{Z}_{[\mathcal{C}_{l}]k}&=\frac{1}{n_{\mathcal{C}_{l}}N}\sum_{i\in\mathcal{C}_{l}}\sum_{m=1}^{M}\sum_{j=1}^{p}(T_{im}(j)-\mathbb{E}[T_{im}(j)])V_{jk}.\end{split}$

### v2

@@ D.1 Deterministic Bounds
$\begin{split}\|\widehat{W}(P^{-1}A-\widehat{A})\|_{F}^{2}&\leq 2\langle\widehat{W}(\widehat{A}-P^{-1}A),(W-\widehat{W}P^{-1})A+Z\rangle\\ &=2\langle\widehat{W}(\widehat{A}-P^{-1}A),(W-\widehat{W}P^{-1})A\rangle+2\langle\widehat{W}(\widehat{A}-P^{-1}A),Z\rangle\\ &\leq 2\|\widehat{W}(\widehat{A}-P^{-1}A)\|_{F}\|(W-\widehat{W}P^{-1})A\|_{F}\\ &\qquad+2\|\widehat{W}(\widehat{A}-P^{-1}A)\|_{F}\max_{U\in\mathbb{R}^{n\times p}:\|U\|_{F}=1}\langle U,Z\rangle\\ \end{split}$
Plugging the upper bound on $\max_{U\in\mathbb{R}^{n\times p}:\|U\|_{F}=1}\langle U,Z\rangle$ which we prove below,
$\begin{split}\|\widehat{W}(P^{-1}A-\widehat{A})\|_{F}&\leq 2\|(W-\widehat{W}P^{-1})A\|_{F}+C_{2}\sqrt{\frac{\log(n)}{N}}\\ &\leq 2\lambda_{1}(A)\|(W-\widehat{W}P^{-1})\|_{F}+C_{2}\sqrt{\frac{\log(n)}{N}}\end{split}$
$\begin{split}&\|P^{-1}A-\widehat{A}\|_{F}\\ &\leq\frac{1}{\lambda_{\min}(\widehat{W})}\left(2\lambda_{1}(A)\|W-\widehat{W}P^{-1}\|_{F}+C_{2}\sqrt{\frac{\log(n)}{N}}\right)\\ &\leq\frac{1}{\lambda_{K}(W)-\|W-\widehat{W}P^{-1}\|_{op}}\left(2\lambda_{1}(A)\|W-\widehat{W}P^{-1}\|_{F}+C_{2}\sqrt{\frac{\log(n)}{N}}\right)\qquad(*)\\ &\leq 2C_{1}\lambda_{1}(A)\|W-\widehat{W}P^{-1}\|_{F}+C_{1}C_{2}\sqrt{\frac{\log(n)}{N}}\end{split}$
@@
Also, assume that $N$ is large enough so that the condition on $N$ in Theorem 2 holds. Combining Lemma 7 and Theorem 3, $\|W-\widehat{W}P^{-1}\|_{F}$ becomes small enough so that $\|W-\widehat{W}P^{-1}\|_{F}<1\leq\lambda_{K}(W)$. Thus, $\frac{1}{\lambda_{K}(W)-\|W-\widehat{W}P^{-1}\|_{F}}\leq C_{1}$ for $C_{1}>1$.
@@
Lastly, we use the fact, $\lambda_{1}(A)\leq\|A\|_{F}\leq\sqrt{K}$ to get the final bound of $A$,
@@
where the error is controlled by the first term, $2C_{1}\lambda_{1}(A)\|W-\widehat{W}P^{-1}\|_{F}$. ∎
Appendix E Auxiliary Lemmas
Let matrices $X,M,Z,W,\text{ and }A$ be defined as (2) in the main manuscript. In this section, we provide inequalities on the singular values of the unobserved quantities $W,M,\text{ and }H$, perturbation bounds for singular spaces, as well as concentration bounds on noise terms, which are useful for proving our main results in Section 3.
E.1 Inequalities on Unobserved Quantities
Lemma 6.
For the matrix $M$, the following inequalities hold:
@@ E.1 Inequalities on Unobserved Quantities
Proof.
We observe that for each $j\in[p]$, the variational characterization of the smallest eigenvalue of the matrix $M^{\top}M/n$ yields:
@@
Lemma 7 (Lemma 6 from the supplemental material of Klopp et al. (2021)).
Let Assumption 2 be satisfied. For the matrices $W$, $H$, $\widehat{H}$ defined in (6) and (7) of the main manuscript, we have
@@
E.2 Matrix Perturbation Bounds
In this section, we provide rate-optimal bounds for the left and right singular subspaces. While the original Wedin’s perturbation bound (Wedin 1972) treats the singular subspaces symmetrically, work by Cai & Zhang (2018) provides sharper bounds for each subspace individually. This refinement is particularly relevant in our setting where an additional denoising step of the left singular subspace leads to different perturbation behaviors of left and right singular subspaces as iterations progress.
@@ E.2 Matrix Perturbation Bounds
Let $X=M+Z$ be a perturbed version of $M$ with $Z$ denoting a perturbation matrix. We can write the SVD of $X$ as:
@@
Lemma 8 (Adapted from Theorem 1 of Cai & Zhang (2018)).
Let $M$, $X$, and $Z$ be as given in Equations (55)-(57). Then:
@@
Proof.
This result is a simplified version of the original theorem under the setting $rank(M)=K$. ∎
E.3 Concentration Bounds
We first introduce the general Bernstein inequality and its variant which will be used for proving high probability bounds for noise terms in Section E.3.2.
E.3.1 General Inequalities
Lemma 9 (Bernstein inequality (Corollary 2.11, Boucheron et al. (2013))).
@@ E.3.1 General Inequalities
Lemma 10 (Bernstein inequality for bounded variables (Theorem 2.8.4, Vershynin (2018))).
@@
E.3.2 Technical Lemmas
Lemma 11 (Concentration of the cross-terms $Z_{i}^{T}Z_{j}$).
Let Assumptions 1-5 hold. With probability at least $1-o(n^{-1})$:
@@ E.3.2 Technical Lemmas
Proof.
The proof is a re-adaptation of Lemma C.4 in Tran et al. (2023) for any word $j$
Similar to the analysis of Ke & Wang (2017), we rewrite each row $X_{i\cdot}$ as a sum over $N$ word entries $T_{im}\in\mathbb{R}^{p}$, where $T_{im}$ denotes the $m^{th}$ word in document $i$, encoded as a one-hot vector:
@@
In the previous expression, under the pLSI model, the $\{T_{im}\}_{m=1}^{N}$ are i.i.d. samples from a multinomial distribution with parameter $M_{i\cdot}$.
@@
so $Q$ is a (double) summation of mutually independent variables. We have $|Q_{im}|\leq 1$, $\mathbb{E}(Q_{im})=0$ and $\mathbb{E}(Q^{2}_{im})\leq{h}_{j}{h}_{l}$ where $\forall j\in[p],h_{j}=\sum_{k=1}^{K}A_{kj}$. The rest of the proof for $V_{2}$ is similar to the standard proof for the usual Bernstein’s inequality.
@@
Lemma 12 (Concentration of the covariance matrix $X^{\top}X$).
@@
$\begin{split}\|Z^{\top}Z-\mathbb{E}[Z^{\top}Z]\|_{F}^{2}&=\sum_{j,j^{\prime}=1}^{p}((Z^{\top}Z)_{jj^{\prime}}-\mathbb{E}[(Z^{\top}Z)_{jj^{\prime}}])^{2}\\ &=\sum_{j}^{p}((Z^{\top}Z)_{jj}-\mathbb{E}[(Z^{\top}Z)_{jj}])^{2}+\sum_{j\neq j^{\prime}}^{p}((Z^{\top}Z)_{jj^{\prime}}-\mathbb{E}[(Z^{\top}Z)_{jj^{\prime}}])^{2}\\ &=\sum_{j}^{p}\left(2(C^{*})^{2}\frac{nh_{j}^{2}\log(n)}{N}+2\frac{(C^{*})^{2}}{N^{2}}\frac{nh_{j}\log(n)}{N}\right)+\sum_{j\neq j^{\prime}}^{p}(C^{*})^{2}\frac{nh_{j}h_{j^{\prime}}\log(n)}{N}\\ &\leq C^{*}\sum_{j,j^{\prime}}^{p}\frac{nh_{j}h_{j^{\prime}}\log(n)}{N}\text{ since by Assumption 5, }\min_{j}h_{j}\geq c_{\min}\frac{\log(n)}{N}\\ &\leq C^{*}K^{2}\frac{n\log(n)}{N}\text{ since }\sum_{j}h_{j}=K\end{split}$
@@
since we assume that $p\ll n$. Therefore, with probability at least $1-o(n^{-1})$:
@@
Noting that $\lambda_{1}(M)\leq\sqrt{n}$ (Lemma 6), and by Lemma 15, with probability at least $1-o(n^{-1})$:
@@
Lemma 13 (Concentration of $\|(\Gamma^{\dagger}Z)_{e\cdot}\|_{2},e\in\mathcal{E}$).
@@
    Each $\eta_{im}$ verifies Bernstein’s condition (60): We have:
@@
    Each of the variance $\mathrm{Var}(S_{m})=\sum_{i=1}^{n}\mathrm{Var}(\eta_{im})$ is also bounded:
@@
with $c_{0}<2$. Therefore, $\mathbb{P}[\exists e\in\mathcal{E}:\quad\|(\Gamma^{\dagger})^{\top}Z)_{e\cdot}\|^{2}\geq C^{*}\rho^{2}(\Gamma){\frac{K\log(n)}{N}}]=o(\frac{1}{n})$ for a choice of $C^{*}$ sufficiently large. ∎
Lemma 14 (Concentration of $\|\Pi Z\|_{F}$).
@@
Proof.
We remind the reader that letting $\mathcal{C}_{j}$ denote the $j^{th}$ connected component of the graph $\mathcal{G}$ and $n_{\mathcal{C}_{l}}=|\mathcal{C}_{l}|$ its cardinality, $\Pi$ can be arranged in a block diagonal form where each block represents a connected component, $\Pi_{[\mathcal{C}_{l}]}=\frac{1}{n_{\mathcal{C}_{l}}}\mathbf{1}_{\mathcal{C}_{l}}\mathbf{1}_{\mathcal{C}_{l}}^{T}$. Since the components $\mathcal{C}_{l}$ are all disjoint, $\|\Pi Z\|_{F}$ can be further decomposed as:
$\begin{split}\|\Pi Z\|_{F}^{2}&=\sum_{l=1}^{n_{C}}\|\frac{1}{n_{\mathcal{C}_{l}}}\mathbf{1}_{\mathcal{C}_{l}}\mathbf{1}_{\mathcal{C}_{l}}^{T}Z_{[\mathcal{C}_{l}]}\|_{F}^{2}\\ &=\sum_{l=1}^{n_{C}}n_{\mathcal{C}_{l}}\|\frac{1}{n_{\mathcal{C}_{l}}}\mathbf{1}_{\mathcal{C}_{l}}^{T}Z_{[\mathcal{C}_{l}]}\|_{2}^{2}\end{split}$
@@
In the previous expression, under the pLSI model, the $\{T_{im}\}_{m=1}^{N}$ are i.i.d. samples from a multinomial distribution with parameter $M_{i\cdot}$. Thus, for each word $j$ and each connected component $\mathcal{C}_{l}$:
$\begin{split}\frac{1}{n_{\mathcal{C}_{l}}}\mathbf{1}_{\mathcal{C}_{l}}^{T}Z_{[\mathcal{C}_{l}]j}&=\frac{1}{n_{\mathcal{C}_{l}}N}\sum_{i\in\mathcal{C}_{l}}\sum_{m=1}^{N}(T_{im}(j)-\mathbb{E}[T_{im}(j)]).\end{split}$
@@
$\begin{split}&\mathbb{P}\left[\exists j\in[p],\exists l\in[n_{C}]:\quad\frac{1}{n_{\mathcal{C}_{l}}N}|\sum_{i\in\mathcal{C}_{l}}\sum_{m=1}^{N}S^{(j)}_{im}|>\sqrt{C^{*}\frac{h_{j}}{n_{\mathcal{C}_{l}}N}\log(n)}\right]\\ &\leq 2pn_{C}\exp\{-{C^{*}\log(n)}\}\\ &=\exp\{\log(2)+\log(p)+\log(n_{C})-C^{*}\log(n)\}\\ &\leq\exp\{-(C^{*}-3)\log(n)\},\end{split}$
As a consequence of Assumption 5, we know that $p\ll n$ (see Remark 2) and under the graph-aligned setting, $n_{\mathcal{C}}\ll n$. Thus with probability $1-o(n^{-1})$, for all $j\in[p]$ and all $l\in[n_{C}]$:
$\begin{split}\|\frac{1}{n_{\mathcal{C}_{l}}}\mathbf{1}_{\mathcal{C}_{l}}^{T}Z_{[n_{C}]}\|_{2}^{2}&\leq\sum_{j\in[p]}C^{*}\frac{h_{j}}{n_{\mathcal{C}_{l}}N}\log(n)=C^{*}\frac{K}{n_{\mathcal{C}_{l}}N}\log(n).\end{split}$
@@
Lemma 15 (Concentration of $\|U^{\top}Z\|_{F}$).
Let Assumptions 1-5 hold. Let $U\in\mathbb{R}^{n\times r}$ denote a projection matrix: $U^{T}U=I_{r}$, with $r$ a term that does not grow with $n$ or $p$ and $r\leq n$, and let $Z$ denote some centered multinomial noise as in $X=M+Z$. Then with probability at least $1-o(n^{-1})$:
@@
Proof.
Let $\tilde{Z}=U^{\top}Z.$ We have:
@@
Thus, ${\tilde{Z}}_{kj}$ is a sum of $N$ centered independent variables.
@@
Since we assume that $pr\ll n$, the result follows. ∎
Lemma 16 (Concentration of $\|\Pi ZV\|_{F}$).
Let Assumptions 1-5 hold. Let $V$ be a orthogonal matrix: $V\in\mathbb{R}^{p\times K},V^{\top}V=I_{K}$. Let $\Pi$ denote the projection matrix unto $\text{Ker}(\Gamma^{\dagger}\Gamma)$, such that $I_{n}=\Pi\oplus^{\perp}\Gamma^{\dagger}\Gamma$. With probability at least $1-o(n^{-1})$:
@@
Proof.
We follow the same procedure as the proof of Lemma 14. Letting $\mathcal{C}_{j}$ denote the $j^{th}$ connected component of the graph $\mathcal{G}$ and $n_{\mathcal{C}_{l}}=|\mathcal{C}_{l}|$ its cardinality, $\|\Pi\tilde{Z}\|_{F}$ can be decomposed as:
@@
$\begin{split}\frac{1}{n_{\mathcal{C}_{l}}}\mathbf{1}_{\mathcal{C}_{l}}^{T}\tilde{Z}_{[\mathcal{C}_{l}]k}&=\frac{1}{n_{\mathcal{C}_{l}}N}\sum_{i\in\mathcal{C}_{l}}\sum_{m=1}^{N}\sum_{j=1}^{p}(T_{im}(j)-\mathbb{E}[T_{im}(j)])V_{jk}.\end{split}$